Cummins India Ltd has the following capital structure, which it considers optimal:

Debt 25%

Preference Shares 10%

Equity shares 65%

Total 100%

Applicable tax rate for the company is 25%. Risk free rate of return is 6%, average equity

market investment has expected rate of return of 12%. The company’s beta is 1.10.

Following terms would apply to new securities being issued as follows:

1. New preference can be issued at a face value of Rs. 100 per share, dividend and cost of

issuance will be Rs. 10 per share and Rs. 2 per share respectively.

2. Debt will bear an interest rate of 9%.

Calculate

a. component cost of debt, preference shares and equity shares assuming that the company

does not issue any additional equity shares.

b. WACC.


1
Expert's answer
2021-05-30T14:16:22-0400

a) Cost of Debt is given to be 9%.

By issuing Preference shares, CP India Ltd can raise a net amount of Rs.100 - Rs.2 = Rs.98 and pay a dividend of Rs.10 annually.

So, cost of preference shares to the company is: Rs.10/Rs.98 = 0.102 or 10.2%.

Cost of equity = risk free rate + beta × (risk premium rate - risk free rate) = 6% + 1.10×(12% - 6%) = 12.6%.

b) WACC = weight of debt × cost of debt × (1 - tax) + Weight of preference shares × cost of preference shares + weight of equity × cost of equity = 0.25×9%×(1 - 0.25) + 0.15×10.2% + 0.60×12.6% = 10.78%.